OpenStudy (anonymous):

help!?!A bird (B) is spotted flying 5,000 feet from a tree (T). An observer (O) spots the bird (B) at a distance of 13,000 feet. What is the angle of depression from the bird (B) to the observer (O)?(will give medal)

OpenStudy (anonymous):

22.70° 44.62° 67.38° 68.96°

OpenStudy (anonymous):

@Directrix can u help

Directrix (directrix):

You will need right triangle trig ratios for this problem

Directrix (directrix):

Which trig ratio is the adjacent side over the hypotenuse?

OpenStudy (anonymous):

cos

Directrix (directrix):

cos(x) = 5000/13000 cos(x) = 5/13 Do you know how to use the inverse cosine button or arc cosine button on your calculator to get the angle x?

Directrix (directrix):

Be sure you have your calculator set for degrees.

OpenStudy (anonymous):

thx